OpenStudy (anonymous):

Find the sum of the given polynomials. ax + by + c, 2ax - 3by + c, and by - c 3ax - by + c 3a + x - bc 3ax + by - cd bx - 3bc + a

OpenStudy (anonymous):

\[ax + by + c + 2ax - 3by + c- c \] is that what you need to find?

OpenStudy (anonymous):

I don't know. I don't understand any of it. I'm just trying to get through this class.

OpenStudy (anonymous):

i think that is it, because it says "find the sum" which means "add"

OpenStudy (anonymous):

Lol I know what "sum" means :p

OpenStudy (anonymous):

\[ax + by + c + 2ax - 3by + c+by- c\]

OpenStudy (anonymous):

ok then we put plus signs between them and combine like terms

OpenStudy (anonymous):

Plus signs in front of everything?

OpenStudy (anonymous):

put plus signs between the polynomials, like i wrote above

OpenStudy (anonymous):

But there's two minus signs

OpenStudy (anonymous):

\[ax + by + c \overbrace{+}^{here} 2ax - 3by + c\overbrace{+}^{here}by- c\]

OpenStudy (anonymous):

that is what i meant by putting plus signs between them

OpenStudy (anonymous):

Yeah the answer I got from that isn't one of the answer choices

OpenStudy (anonymous):

\(ax + by + c\) plus \( 2ax - 3by + c\) plus \(by - c \) ok lets combine the like terms, meaning the terms with \(x\), the terms with \(y\) and the constant \(c\) we can cheat on this one because for the \(c\) terms you have \(+c+c-c\)

OpenStudy (anonymous):

and \(c+c-c=c\) just like \(5+5-5=5\) only one answer has \(c\) by itself, so although we can do the rest as well, that has to be the right one

OpenStudy (anonymous):

i am not sure what you got, but for the \(x\) terms you have \(ax+2ax\) and that gives you \(3ax\)

OpenStudy (anonymous):

you also have \(by-3by+by\) and since \(1-3+1=-1\) you have a total of \(-by\)

OpenStudy (anonymous):

final answer\[3ax-by+c\]

OpenStudy (anonymous):

so it looks like, for a change, the first answer is right not sure if this made any sense, but it is the right one